OpenStudy (anonymous):

if f(x)=-9x-2 and g(x)=sqrt{x-5} , which statement is true -2 is in the domain of f o g -2 is not in the domain of fog ???

OpenStudy (anonymous):

f o g *

OpenStudy (fifciol):

f o g =f(g(x))=f(sqrt{x-5})=-9(sqrt{x-5})-2) the domain of this function : \[x-5\ge 0\] \[x \ge 5\] so x=-2 2 is not in the domain of f o g

OpenStudy (fifciol):

x=-2 is not in the domain of fog *

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

wanna help me with some more ?

OpenStudy (fifciol):

of course :)